OpenStudy (anonymous):

Hey guys, i need help to simplify this log: log(1/2) (1/8)^(2-x) (1/2) is the log base.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the answer is 1/2x

OpenStudy (anonymous):

\[\log_{1/2}\left(\frac{1}{8}\right)^{2-x}\]

OpenStudy (anonymous):

that the question?

OpenStudy (anonymous):

pull the \(2-x\) out front as \[(2-x)\log_{1/2}\left(\frac{1}{8}\right)\] first

OpenStudy (anonymous):

then figure out what \[\log_{1/2}\left(\frac{1}{8}\right)\] is, which should not bet too hard

OpenStudy (anonymous):

solve \[\left(\frac{1}{2}\right)^y=\frac{1}{8}\] for \(y\) in your head and you will get it

OpenStudy (anonymous):

thanks for replying, i just need to simplify, not find the x the answer in the test says 6-3x